Show Solution.

Sphere A (8q) touched Sphere B (12q) then separated. Sphere A touched Sphere C (-5q) then separated.

a) What is the Σq on the three spheres before touching each other?

b) What are the final charges on each sphere, assuming they are identical in radius, which is r?


1
Expert's answer
2021-02-09T15:26:04-0500

(a)

"Q_{tot}=Q_A+Q_B+Q_C=8q+12q-5q=15q."

(b) When sphere A is touched to sphere B, the net charge on each of them equals:


"Q=\dfrac{Q_A+Q_B}{2}=\dfrac{8q+12q}{2}=10q."

Then, sphere A is touched to sphere C, so the net charge on each of them equals:


"Q=\dfrac{Q_A+Q_C}{2}=\dfrac{10q-5q}{2}=2.5q."

Therefore, each sphere has the following final charges:


"Q_A=2.5q, Q_B=10q, Q_C=2.5q."